# LATex: Leveraging Attribute-based Text Knowledge for Aerial-Ground Person Re-Identification

### v1

@@ Abstract
Aerial-Ground person Re-IDentification (AG-ReID) aims to retrieve specific persons across heterogeneous cameras in different views. Previous methods usually adopt large-scale models, focusing on view-invariant features. However, they overlook the semantic information in person attributes. Additionally, existing training strategies often rely on full fine-tuning large-scale models, which significantly increases training costs. To address these issues, we propose a novel framework named LATex for AG-ReID, which adopts prompt-tuning strategies to Leverage Attribute-based Text knowledge. More specifically, we first introduce the Contrastive Language-Image Pre-training (CLIP) model as the backbone, and propose an Attribute-aware Image Encoder (AIE) to extract global semantic features and attribute-aware features. Then, with these features, we propose a Prompted Attribute Classifier Group (PACG) to generate person attribute predictions and obtain the encoded representations of predicted attributes. Finally, we design a Coupled Prompt Template (CPT) to transform attribute tokens and view information into structured sentences. These sentences are processed by the text encoder of CLIP to generate more discriminative features. As a result, our framework can fully leverage attribute-based text knowledge to improve the AG-ReID. Extensive experiments on three AG-ReID benchmarks demonstrate the effectiveness of our proposed LATex. The source code will be available.
1 Introduction
Person Re-IDentification (ReID) aims to retrieve the same individual across different cameras (Jiang et al., 2024; Zheng et al., 2016). In recent years, ReID has gained significant attention (Lu et al., 2023; Shi et al., 2023, 2024; Yin et al., 2024; Gao et al., 2024; Liu et al., 2023; Zhang et al., 2021) due to its wide range of applications, including video surveillance and public security. More recently, ReID across heterogeneous camera views, especially Aerial-Ground person ReID (AG-ReID), has become a more realistic scenario (Nguyen et al., 2023a, b; Zhang et al., 2024b) due to the development of drones and advancements in aerial surveillance. Different from traditional ReID tasks (Crawford et al., 2023; Leng et al., 2019; Li et al., 2017), AG-ReID amplifies the challenges posed by viewpoint variations due to drastic changes between different cameras. These variations significantly affect the distribution of body parts (Schumann & Metzler, 2017), making it more difficult to learn robust features that remain consistent across diverse perspectives.
Figure 1: An example of a person captured by an aerial camera (a) and a ground camera (c), along with the corresponding attributes (b). Despite significant visual variations in the images caused by viewpoint changes, person attributes remain consistent.
@@ 1 Introduction
Previous methods (Nguyen et al., 2023a; Zhang et al., 2024b) focus on mitigating the negative effects of drastic perspective changes and learning viewpoint-robust image features. However, they often overlook the potential of leveraging person attribute information. As shown in Fig. 1(a) and Fig. 1(c) , different camera viewpoints result in significant visual differences. Despite these significant viewpoint changes, person attributes such as ethnicity, gender and clothing remain unaffected. This stability provides consistent information to obtain robust cross-view features. Meanwhile, existing methods (Nguyen et al., 2023a; Zhang et al., 2024b) rely on the full fine-tuning strategy, which significantly increases the training cost. Fortunately, prompt tuning (Jia et al., 2022; Han et al., 2024; Lester et al., 2021; Liu et al., 2021a) offers a way to reduce the training cost while effectively integrating the pre-trained knowledge of large-scale models like CLIP (Radford et al., 2021) into specific domains (Li et al., 2023; Wang et al., 2024a). This approach has proven to be more efficient since it reduces the number of trainable parameters and lowers the training cost, while still leveraging the pre-trained model’s powerful knowledge. Motivated by these observations, we propose a novel framework named LATex for AG-ReID, which leverages attribute-based text knowledge to enhance the feature discrimination with the prompt tuning strategies.
@@
Specifically, our framework consists of three key components: Attribute-aware Image Encoder (AIE), Prompted Attribute Classifier Group (PACG) and Coupled Prompt Template (CPT). First, we introduce AIE to fine-tune CLIP with learnable prompts, transferring the powerful pre-trained knowledge to AG-ReID. In addition, AIE incorporates attribute tokens to enable fine-grained perception of person attribute information. Then, PACG is employed to further enhance AIE’s attribute perception capabilities and to generate attribute-based text tokens. Afterwards, CPT is proposed to integrate attribute-based tokens and view-informed tokens into structured sentences. Finally, these sentences are processed by CLIP’s text encoder, enabling more accurate person retrieval across different camera views by explicitly leveraging person attributes. Extensive experiments on three AG-ReID benchmarks fully validate the effectiveness and efficiency of our proposed framework.
@@
  • New insight. We observe the distinct benefits of person attribute information for AG-ReID tasks. This insight prompts us to consider the problem from an attribute-based perspective. Furthermore, we introduce a practical new method to mitigate the challenges in target retrieval posed by drastic viewpoint changes.
  • Novel framework. We present LATex, a novel feature learning framework for AG-ReID that leverages attribute-based text knowledge with prompt-tuning. It not only reduces the training cost, but also extracts more discriminative features.
  • Effective modules. We propose two effective modules, i.e., PACG and CPT. PACG can effectively predict person attributes and generate embedded tokens of predicted attributes. CPT integrates text knowledge by transforming attribute tokens and view information into structured sentences.
@@
2 Related Works
Person ReID is a longstanding task in computer vision and machine learning, gaining significant attention due to its wide range of real-world applications. Previous research has primarily focused on view-homogeneous scenarios, where all cameras in the network are assumed to operate under the same viewpoint. In this setting, notable advancements have been achieved, primarily categorized into CNN-based methods, such as PCB (Sun et al., 2021), and Transformer-based methods, such as TransReID (He et al., 2021). For multi-modal ReID, more advanced approaches include EDITOR (Zhang et al., 2024a), DeMo (Wang et al., 2024d), MambaPro (Wang et al., 2024b), TOP-ReID (Wang et al., 2024c). However, these methods rely on well-deployed camera networks and are significantly limited in sparsely covered areas due to the lack of candidate images.
@@ 2 Related Works
Fortunately, advancements in Unmanned Aerial Vehicle (UAV) technology have made it feasible to deploy cameras on UAV, enhancing surveillance coverage in regions with sparse ground camera networks. However, the substantial viewpoint variations between UAV cameras and fixed ground cameras pose significant challenges, as directly transferring previous view-homogeneous methods often leads to suboptimal performance. To address this issue, Nguyen et al. (Nguyen et al., 2023b, 2024) propose using person attributes as soft labels for supervision, while VDT (Zhang et al., 2024b) attempts to separate identity-related features from viewpoint-specific features by employing view tokens and an orthogonal loss. Despite effective, these methods ignore the benefits of explicitly using attributes for cross-view retrieval. In this work, we propose LATex, an framework that leverages attribute information to achieve outstanding performance with significantly fewer trainable parameters.
Figure 2: The illustration of our idealized assumption and results. Under ideal conditions, with the assistance of person attributes, the model achieves nearly perfect scores on validation metrics.
@@
Figure 3: The illustration of the proposed LATex framework. The Attribute-aware Image Encoder (AIE) first extracts global semantic features and attribute-aware features. Then, the Prompted Attribute Classifier Group (PACG) generates person attribute predictions and obtain embedded tokens of predicted attributes. Afterwards, the Coupled Prompt Template (CPT) transforms attribute tokens and view information into structured sentences. Finally, the structured sentences are processed by the text encoder of CLIP to generate discriminative features for person retrieval.
3 Preliminaries
3.1 Problem Definition
@@
We focus on the task where each person is captured from different camera platforms, such as CCTV or UAV. Our goal is to enable the model to correctly match the query image with the gallery image. Formally, we define the problem as follows: Given a training dataset $\mathcal{C}=(\mathcal{C}^{Img},\mathcal{C}^{ID},\mathcal{C}^{View})$, where $\mathcal{C}^{Img}=\{\mathcal{C}^{Img}_{1},\cdots,\mathcal{C}^{Img}_{B}\}$ is a set with $B$ images. $\mathcal{C}^{ID}\in\mathbb{R}^{B}$ and $\mathcal{C}^{View}\in\mathbb{R}^{B}$ denote the corresponding IDentity (ID) labels and view labels, respectively. We consider a deep model $\mathcal{M}$ with trainable parameters $\theta$ to extract discriminative person representations from input images:
@@
where $F_{i}$ and $F_{j}$ are the representations extracted by the deep model $\mathcal{M}$, respectively. Here, the input data consists of any two instances, i.e., $(\mathcal{C}^{Img}_{i},\mathcal{C}^{ID}_{i},\mathcal{C}^{View}_{i})$ and $(\mathcal{C}^{Img}_{j},\mathcal{C}^{ID}_{j},\mathcal{C}^{View}_{j})$. Our training objective is to optimize the trainable parameters $\theta$ such that during the inference phase, the following condition holds:
$\begin{cases}\mathcal{D}_{pos}=\mathcal{D}(F_{i},F_{j})&\text{if }\mathcal{C}^{ID}_{i}=\mathcal{C}^{ID}_{j},\\ \mathcal{D}_{neg}=\mathcal{D}(F_{i},F_{j})&\text{if }\mathcal{C}^{ID}_{i}\neq\mathcal{C}^{ID}_{j},\\ \mathcal{D}_{pos}\ll\mathcal{D}_{neg},\\ \end{cases}$ (3)
@@
where $\mathcal{D}(\cdot)$ denotes a certain distance metric for person representations. Given a query person, we use this distance to measure, rank, and subsequently match the gallery image corresponding to the same person identity.
3.2 Motivation and Idealized Assumption
Motivation. Previous methods have shown strong performance for person ReID tasks with the same viewpoints but struggle in scenarios with significant viewpoint variations. Based on our daily experiences and observations of the dataset (Fig.1), we notice that person attributes remain robust in such complex scenarios, providing valuable additional information for target person retrieval. In light of this, we try to fully exploit this advantage to alleviate the challenges posed by viewpoint changes in ReID tasks.
Idealized Assumption. We consider an ideal scenario where the ground truth of all person attributes is known in both the training and testing sets. Under this premise, we construct a simple idealized assumption using prompt-tuning CLIP to validate the feasibility of our motivation. This setup represents the theoretical upper bound of our framework’s performance. Fig. 2 (a) provides an overview of our idealized assumption, with additional details available in the appendix. And Fig. 2 (b) shows the results of our idealized assumption on the AG-ReID.v1 dataset, where our method achieves near-perfect scores on both protocols, fully demonstrating the validity of our motivation. Despite the superior performance, we recognize that the above experimental setting is overly idealized and impractical in real-world applications. Specifically, our proposed idealized assumption relies on the strong assumption that all person attributes are known during inference. However, annotating attributes for the vast amounts of data collected by numerous surveillance cameras is infeasible in practice. To address this, we propose a feature learning framework, which enables the model to automatically generate person attributes during inference. This eliminates the aforementioned unrealistic assumption, significantly enhancing the feasibility of our approach.
4 Proposed Method
4.1 Attribute-aware Image Encoder
@@
To transfer the rich knowledge of CLIP (Radford et al., 2021) to the ReID task and extract attribute-specific information, we propose the Attribute-aware Image Encoder (AIE). Previous Transformer-based approaches (Nguyen et al., 2023a; Zhang et al., 2024b) typically rely on full fine-tuning, which leads to very high training costs. To address this issue, we adopt prompt-tuning strategies to extract discriminative features. Formally, given the input images $\mathcal{V}\in\mathbb{R}^{H\times W\times 3}$ from different views, we embed $\mathcal{V}$ to obtain the visual embedding $F^{0}=[F_{v}^{0},f_{v}^{0}]$, where $F_{v}^{0}\in\mathbb{R}^{C}$ is the class token and $f_{v}^{0}\in\mathbb{R}^{N\times C}$ is the patch tokens. Here, $C$ is the dimension of the token embeddings while $N$ is the total number of patches. Then, for the $i$-th Transformer layer $\varOmega_{i}$, we denote $P^{i}_{v}\in\mathbb{R}^{T\times C}$ as the learnable prompts, i.e., $P^{i}_{v}=\{P^{i,1}_{v},\cdots,P^{i,T}_{v}\}$. Here, $T$ is the total number of person attributes. The learnable prompts $P^{i}_{v}$ are concatenated with $F^{i}$, enabling the perception of attribute-specific information, and then passed through $\varOmega_{i}$ as follows:
$[F^{i+1},\tilde{P}^{i+1}_{v}]=\varOmega_{i}([F^{i},P^{i}_{v}]).$ (4)
@@
Here, [$\cdot$] means the concatenation operation along the token dimension. Finally, we extract the attribute prompts $\tilde{P}^{L}_{v}\in\mathbb{R}^{T\times C}$ and the visual class token $F_{v}^{L}\in\mathbb{R}^{C}$ from the final layer of AIE for further processing.
4.2 Prompted Attribute Classifier Group

### v2

@@ Abstract
As an important task in intelligent transportation systems, Aerial-Ground person Re-IDentification (AG-ReID) aims to retrieve specific persons across heterogeneous cameras in different viewpoints. Previous methods typically adopt deep learning-based models, focusing on extracting view-invariant features. However, they usually overlook the semantic information in person attributes. In addition, existing training strategies often rely on full fine-tuning large-scale models, which significantly increases training costs. To address these issues, we propose a novel framework named LATex for AG-ReID, which adopts prompt-tuning strategies to leverage attribute-based text knowledge. More specifically, we first introduce the Contrastive Language-Image Pre-training (CLIP) model as the backbone, and propose an Attribute-aware Image Encoder (AIE) to extract both global semantic features and attribute-aware features from input images. Then, with these features, we propose a Prompted Attribute Classifier Group (PACG) to predict person attributes and obtain attribute representations. Finally, we design a Coupled Prompt Template (CPT) to transform attribute representations and view information into structured sentences. These sentences are processed by the text encoder of CLIP to generate more discriminative features. As a result, our framework can fully leverage attribute-based text knowledge to improve AG-ReID performance. Extensive experiments on three AG-ReID benchmarks demonstrate the effectiveness of our proposed methods. The source code will be available.
Index Terms:
Aerial-Ground Person Re-identification, Image-Text Retrieval, Attribute Prediction, Prompt Learning.
I Introduction
@@
Person Re-IDentification (ReID) aims to retrieve the same individual across different cameras. In recent years, ReID has attracted considerable interest [1, 2, 3, 4, 5, 6, 7, 8] due to its wide range of applications, including intelligent surveillance and transportation system. More recently, ReID across heterogeneous camera viewpoints, especially Aerial-Ground person ReID (AG-ReID), has become a more realistic transportation scenario [9, 10, 11] due to the development of drones and advancements in aerial surveillance. Unlike traditional ReID tasks [12, 13, 14], AG-ReID amplifies the challenges posed by viewpoint variations due to drastic changes between different cameras. These variations significantly affect the distribution of body parts [15], making it more difficult to learn robust features that remain consistent across diverse views. Therefore, previous methods [9, 11] focus on mitigating the negative effects of drastic view changes and learning viewpoint-robust image features. However, they often overlook the potential of leveraging person attribute information. As shown in Fig. 1(a) and Fig. 1(c), different camera viewpoints may result in significant visual differences. Despite these significant visual differences, person attributes such as ethnicity, gender, and clothing remain unaffected. This stability provides consistent information to obtain robust cross-view features. Meanwhile, existing methods [9, 11] rely on the full fine-tuning strategy, significantly raising training costs. Fortunately, prompt-tuning [16, 17, 18, 19] offers a way to reduce the training cost. It also effectively integrates the pre-trained knowledge of large-scale models into specific domains [20, 21].
Figure 1: An example that a person captured under (a) the aerial view by UAV and (c) the ground view by CCTV, along with (b) the corresponding person attributes. Despite significant variations in the images caused by drastic viewpoint changes, person attributes remain consistent.
@@
Motivated by the above observations, we propose a novel framework named LATex for AG-ReID, which leverages attribute-based text knowledge via prompt-tuning strategies to enhance the feature discrimination. Specifically, our framework consists of three key components: an Attribute-aware Image Encoder (AIE), a Prompted Attribute Classifier Group (PACG), and a Coupled Prompt Template (CPT). First, we introduce AIE to fine-tune the Contrastive Language-Image Pre-training (CLIP) model [22] with learnable prompts, transferring the powerful pre-trained knowledge to AG-ReID. In addition, AIE incorporates attribute tokens to enable fine-grained perception of person attribute information. Then, PACG is employed to further enhance AIE’s attribute perception capabilities and generate attribute representations. Afterwards, CPT is proposed to transform attribute representations and view information into structured sentences. Finally, these sentences are processed by CLIP’s text encoder, enabling more accurate person retrieval across different camera viewpoints by explicitly leveraging information hidden in the attributes. Extensive experiments on three AG-ReID benchmarks fully validate the effectiveness of our proposed framework.
@@
  • New insight. We observe the distinct benefits of person attribute information for AG-ReID tasks. This insight inspires us to consider the problem from an attribute-based perspective. Based on the attribute consistency, we introduce a practical method to mitigate the challenges in AG-ReID posed by drastic viewpoint changes.
  • Novel framework. We present LATex, a novel feature learning framework that leverages attribute-based text knowledge with prompt-tuning strategies. It not only reduces the resource requirement during training, but also extracts more discriminative features for AG-ReID.
  • Effective modules. We propose two effective modules, i.e., PACG and CPT. PACG can effectively predict person attributes and generate attribute representations. CPT integrates text knowledge by transforming attribute representations and view information into structured sentences.
@@
II Related Works
II-A View-Homogeneous ReID
Person ReID is a long-standing task in computer vision and machine learning, gaining significant attention due to its wide range of real-world applications [23, 24, 25, 26, 27]. Previous research has primarily focused on view-homogeneous scenarios, where all cameras in the surveillance network are assumed to operate under the similar viewpoint. Coarsely, the view-homogeneous ReID can be categorized into two types: ground-view and aerial-view. In fact, the ground-view ReID has been widely researched with the support of various datasets, such as Market1501 [28], MSMT17 [29] and CUHK03 [30]. As a consequence, notable advancements have been achieved, primarily categorized into CNN-based methods [31, 32, 33] and Transformer-based methods [34, 35, 36]. For CNN-based methods, Sun et al. [37] and Wang et al. [38] enhance global feature representations by dividing the person image into several parts and extracting the part-level features. Furthermore, Luo et al. [39] provide a strong ReID baseline by introducing some useful tricks. Focusing on the computational efficiency, Quan et al. [40] successfully construct a compact model, namely Auto-ReID, to obtain local discriminative features. In recent years, many methods based on Vision Transformer (ViT) have emerged in the ReID community. For example, He et al. [41] first introduce Transformer into person ReID, achieving promising results. Afterwards, many researchers further leverage Transformers to extract more discriminative person representations [42, 7, 36, 43]. Beside the spatial cues, Li et al. [44] extract high-frequency information of person images to obtain stable representations for ReID. In terms of aerial-view ReID, UAV-Human [45] and PRAI-1581 [46] are the primary benchmarks in this person ReID sub-task. As for advanced methods, Qiu et al. [47] introduce a key-point disentangling strategy for aerial-view ReID. To address the challenge of person rotation, Wang et al. [48] propose a rotation exploration for aerial-view ReID. Although effective, these approaches perform poorly under drastic viewpoint changes, which inevitably appear in AG-ReID.
II-B Aerial-Ground ReID
@@
Recently, advancements in Unmanned Aerial Vehicle (UAV) technologies have made it feasible to deploy dynamic cameras, enhancing surveillance coverage in regions with sparse ground camera networks. However, it poses significant challenges due to the substantial viewpoint variations between UAV cameras and fixed ground cameras. As a result, directly transferring previous view-homogeneous ReID methods often leads to suboptimal performance. To address this issue, AG-ReID has been proposed as a new sub-task of person ReID. To achieve the model training and evaluation, Nguyen et al. [10] first collect an outdoor scene dataset with person attribute annotations, namely AG-ReID.v1. Afterwards, they extend the AG-ReID.v1 with more identities and viewpoints, as AG-ReID.v2 [49]. Zhang et al. [11] construct a large-scale synthesized AG-ReID benchmark, named CARGO. Recently, Zhang et al. [50] consider video-based AG-ReID and contribute the first benchmark, named G2A-VReID. Then, Nguyen et al. [51] further contribute a large-scale video-based AG-ReID benchmark, named AG-VPReID. As for technical methods, Nguyen et al. [10, 49] propose multi-stream frameworks and use person attributes as auxiliary labels for supervision. Based on ViT, Zhang et al. [11] separate identity-related features from viewpoint-specific features by employing view tokens and an orthogonal loss. Moreover, Wang et al. [52] consider the person local features and introduce a prompt-based framework for better AG-ReID. On the other hand, Wang et al. [53] employ a dynamic token selection strategy to focus on key person regions. Hu et al. [54] propose a view-aware stable diffusion to extract comprehensive features for AG-ReID. Although effective, these methods ignore the benefits of explicitly using person attributes for cross-view retrieval. Based on the observations in Fig. 1, we notice that person attributes remain robust in complex scenarios, providing valuable information for discriminative features. Motivated by this, we fully exploit this advantage to alleviate the viewpoint changes in AG-ReID tasks. Technically, we propose a new feature learning framework named LATex that predicts and leverages attribute knowledge to achieve better performance with fewer trainable parameters.
II-C Prompt-Tuning in Person ReID
Prompt-tuning aims to transfer the knowledge of pre-trained models to unseen domains via trainable prompts. Moreover, it typically requires fewer computation resources than full fine-tuning, while also achieving superior performance. This property makes prompt-tuning widely applicable across various tasks [16, 20, 18]. In the ReID field, Li et al. [20] first exploit vision-language models with prompt-tuning to address the lack of missing concrete text labels in image-based person Re-ID. Li et al. [55] propose prototypical contrastive learning to fine-tune CLIP for object ReID. Wu et al. [56] enhance visible-infrared person ReID with modality-aware and instance-aware visual prompt learning. Wang et al. [57] introduce diverse prompt-tuning methods to distill CLIP for learning discriminative person shape representations. Li et al. [58] propose person prompts and clothes prompts to learn cloth-agnostic features for cloth-changing person ReID. Very recently, Wang et al. [52] introduce self-calibrating and adaptive prompts for AG-ReID. Different from previous works, we not only employ learnable prompts as additional tokens to help pre-trained large models generalize to ReID domains, but also leverage the prompt knowledge to further enhance discriminative representations for AG-ReID.
@@ II-C Prompt-Tuning in Person ReID
Figure 2: The illustration of the proposed LATex framework. The Attribute-aware Image Encoder (AIE) first extracts global semantic features and attribute-aware features. Then, the Prompted Attribute Classifier Group (PACG) generates person attribute predictions and obtain specific representations of predicted attributes. Afterwards, the Coupled Prompt Template (CPT) transforms attribute representations and view information into structured sentences. Finally, the structured sentences are processed by the text encoder of CLIP to generate discriminative features for person ReID, integrated with global semantic features.
III Proposed Method
III-A Problem Definition
@@
We focus on the AG-ReID task where each person may be captured from different camera platforms, such as CCTV or UAV. Our goal is to enable the model to correctly match the query image with the gallery image. Formally, we define the problem as follows: Given a training dataset $\mathcal{C}=(\mathcal{C}^{Img},\mathcal{C}^{ID},\mathcal{C}^{View})$, $\mathcal{C}^{Img}$, $\mathcal{C}^{ID}$ and $\mathcal{C}^{View}$ denote the number of person images, identity labels and view labels, respectively. We consider a model $\mathcal{M}$ with trainable parameters $\theta$ to extract discriminative representations from input images:
@@
where $F_{i}$ and $F_{j}$ are the representations extracted by the model $\mathcal{M}$, respectively. Here, the input data consists of any two instances, i.e., $(\mathcal{C}^{Img}_{i},\mathcal{C}^{ID}_{i},\mathcal{C}^{View}_{i})$ and $(\mathcal{C}^{Img}_{j},\mathcal{C}^{ID}_{j},\mathcal{C}^{View}_{j})$, where $i\neq j$. Our training objective is to optimize the trainable parameters $\theta$ such that during the inference phase, the following condition holds:
[…]
@@
where $\mathcal{D}(\cdot)$ denotes a certain distance metric. Given a query person image, we use this distance to match the gallery image corresponding to the same person identity.
III-B Attribute-aware Image Encoder
@@
To transfer the rich knowledge of CLIP to the ReID task and extract attribute information, we propose the Attribute-aware Image Encoder (AIE). Previous Transformer-based approaches [9, 11] typically rely on full fine-tuning strategies, which lead to very high training costs. To address this issue, we adopt prompt-tuning strategies to extract discriminative features with reduced training resource requirements. Formally, given the input image $\mathcal{V}\in\mathbb{R}^{H\times W\times 3}$ from different views, we embed $\mathcal{V}$ to obtain the visual embedding $F^{0}=[F_{v}^{0},f_{v}^{0}]$, where $F_{v}^{0}\in\mathbb{R}^{C}$ is the class token and $f_{v}^{0}\in\mathbb{R}^{N\times C}$ are patch tokens. Here, $C$ is the dimension of the token embeddings while $N$ is the total number of patches. Then, for the $i$-th Transformer layer $\varOmega_{i}$, we denote $P^{i}_{v}\in\mathbb{R}^{T\times C}$ as the learnable prompts, i.e., $P^{i}_{v}=\{P^{i,1}_{v},\cdots,P^{i,T}_{v}\}$. Here, $T$ is the total number of person attributes. The learnable prompts $P^{i}_{v}$ are concatenated with $F^{i}$, enabling the perception of attribute-specific information, and then passed through $\varOmega_{i}$ as follows:
$F^{i+1},\tilde{P}^{i+1}_{v}]=\varOmega_{i}([F^{i},P^{i}_{v}]).$ (4)
@@
Here, $\cdot$] means the concatenation operation along the token dimension. Finally, we extract the attribute prompts $\tilde{P}^{L}_{v}\in\mathbb{R}^{T\times C}$ and the visual class token $F_{v}^{L}\in\mathbb{R}^{C}$ from the final layer of AIE for further processing.
III-C Prompted Attribute Classifier Group